### v2

# Automorphism group schemes and Weyl groups of gradings

Abstract.
A scheme theoretic version of the automorphism group of a grading on an algebra is presented, and the classical result that shows that, over algebraically closed fields of characteristic 0, the automorphism group of a grading is the normalizer of its diagonal group is extended, over arbitrary fields, to this scheme setting.
The quotient of the scheme theoretic versions of the automorphism group and the stabilizer of a grading turns out to be a constant group scheme, called the Weyl group scheme of the grading. For algebraically closed fields this is the constant group scheme associated to the ordinary Weyl group of the grading, but this fails over arbitrary fields.
Key words and phrases:
Grading, diagonal group, stabilizer, automorphism group, normalizer, Weyl group.
2020 Mathematics Subject Classification:
Primary 17A36; Secondary 17B70
Supported by grant PID2021-123461NB-C21, funded by MCIN/AEI/10.13039/ 501100011033 and by “ERDF A way of making Europe”; and by grant E22_20R (Gobierno de Aragón).
Dedicated to professor Leonid Makar-Limanov, on the occasion of his 80th birthday
1. Introduction
Let $\mathcal{A}$ be a nonassociative (i.e., not necessarily associative) algebra over a field $\mathbb{F}$ and let $G$ be an abelian group. A grading on $\mathcal{A}$ by $G$, or $G$-grading, is a vector space decomposition
$\Gamma:\mathcal{A}=\bigoplus_{g\in G}\mathcal{A}_{g},$ (1.1)
satisfying $\mathcal{A}_{g}\mathcal{A}_{h}\subseteq\mathcal{A}_{gh}$ for all $g,h\in G$. In this case, the nonzero elements in $\mathcal{A}_{g}$ are said to be homogeneous of degree $g$ and we write $\deg x=g$ for $0\neq x\in\mathcal{A}_{g}$. The subspace $\mathcal{A}_{g}$ is the homogeneous component of degree $g$.
The set
$\operatorname{\mathrm{Supp}}(\Gamma):=\{g\in G\mid\mathcal{A}_{g}\neq 0\}$
is called the support of $\Gamma$.
There are some natural subgroups of the group $\operatorname{\mathrm{Aut}}(\mathcal{A})$ of automorphisms of $\mathcal{A}$ attached to $\Gamma$ (see [PZ89] and [EK13, Chapter 1]). The first one is the stabilizer of the grading:
$\operatorname{\mathrm{Stab}}(\Gamma):=\{\varphi\in\operatorname{\mathrm{Aut}}(\mathcal{A})\mid\varphi(\mathcal{A}_{g})\subseteq\mathcal{A}_{g}\ \forall g\in G\}.$
Another one is the automorphism group of the grading:
$\operatorname{\mathrm{Aut}}(\Gamma):=\{\varphi\in\operatorname{\mathrm{Aut}}(\mathcal{A})\mid\forall g\in G\ \exists h\in G\text{ such that }\varphi(\mathcal{A}_{g})\subseteq\mathcal{A}_{h}\}.$
Each $\varphi\in\operatorname{\mathrm{Aut}}(\Gamma)$ induces a permutation of $\operatorname{\mathrm{Supp}}(\Gamma)$, and this gives an exact sequence
[…]
The quotient
$W(\Gamma):=\operatorname{\mathrm{Aut}}(\Gamma)/\operatorname{\mathrm{Stab}}(\Gamma)$
is called the Weyl group of the grading and is (isomorphic to) a subgroup of the symmetric group $\operatorname{\mathrm{Sym}}\bigl(\operatorname{\mathrm{Supp}}(\Gamma)\bigr)$. The Weyl group measures how much symmetric a grading is.
Assume for a while that the ground field $\mathbb{F}$ is algebraically closed of characteristic 0, and that $\mathcal{A}$ is finite dimensional. Assume too, without loss of generality, that $G$ is finitely generated, otherwise substitute $G$ by the subgroup generated by $\operatorname{\mathrm{Supp}}(\Gamma)$. Consider the group of characters $\widehat{G}=\operatorname{\mathrm{Hom}}(G,\mathbb{F}^{\times})$ (group homomorphisms). The grading $\Gamma$ induces a homomorphism (of algebraic groups)
$\eta_{\Gamma}\colon\widehat{G}\longrightarrow\operatorname{\mathrm{Aut}}(\mathcal{A}),$ (1.2)
that takes any character $\chi\in\widehat{G}$ to the automorphism $\tau_{\chi}\in\operatorname{\mathrm{Aut}}(\mathcal{A})$ given by:
$\tau_{\chi}\colon x\in\mathcal{A}_{g}\longmapsto\chi(g)x\quad\forall g\in G.$ (1.3)
Since characters separate elements of $G$, the homogeneous components of $\Gamma$ can be retrieved from $\eta_{\Gamma}$: for all $g\in G$ we have
$\mathcal{A}_{g}=\{x\in\mathcal{A}\mid\tau_{\chi}(x)=\chi(g)x\ \forall\chi\in\widehat{G}\}$.
The image of $\eta_{\Gamma}$ is contained in the diagonal group of $\Gamma$:
$\operatorname{\mathrm{Diag}}(\Gamma):=\{\varphi\in\operatorname{\mathrm{Stab}}(\Gamma)\mid\forall g\in G\ \exists\mu_{g}\in\mathbb{F}^{\times}\text{ such that }\varphi|_{\mathcal{A}_{g}}=\mu_{g}\mathrm{id}\}.$ (1.4)
One obtains easily the following equalities (see [EK13, (1.6),(1.7)]):
$\operatorname{\mathrm{Stab}}(\Gamma)=\operatorname{\mathrm{Cent}}_{\operatorname{\mathrm{Aut}}(\mathcal{A})}\bigl(\operatorname{\mathrm{Diag}}(\Gamma)\bigr),\quad\operatorname{\mathrm{Aut}}(\Gamma)=\operatorname{\mathrm{Norm}}_{\operatorname{\mathrm{Aut}}(\mathcal{A})}\bigl(\operatorname{\mathrm{Diag}}(\Gamma)\bigr).$ (1.5)
That is the stabilizer (respectively, automorphism group) of $\Gamma$ is the centralizer (resp., normalizer) in $\operatorname{\mathrm{Aut}}(\mathcal{A})$ of the diagonal group.
For arbitrary $\mathbb{F}$, and keeping our hypothesis of finite dimensionality of $\mathcal{A}$, the group of automorphisms $\operatorname{\mathrm{Aut}}(\mathcal{A})$ must be substituted by the affine group scheme of automorphisms $\mathbf{Aut}(\mathcal{A})$.
The functorial point of view of affine group schemes, as considered in [Wat79], [KMRT98, Chapter 6], or [Mil17], will be followed. The reader is referred to these references for their basic facts.
Therefore, an affine group scheme is a representable functor from the category $\operatorname{\mathrm{Alg_{\mathbb{F}}}}$ of unital commutative associative algebras over $\mathbb{F}$ to the category $\operatorname{\mathrm{Grp}}$ of groups. In particular, $\mathbf{Aut}(\mathcal{A})$ acts as follows on objects:
$\mathbf{Aut}(\mathcal{A})\colon R\mapsto\operatorname{\mathrm{Aut}}_{\operatorname{\mathrm{Alg_{R}}}}(\mathcal{A}\otimes_{\mathbb{F}}R)$
(group of automorphisms of $\mathcal{A}\otimes_{\mathbb{F}}R$ as an algebra over $R$). The action on homomorphisms is the natural one by ‘scalar extension’.
Given our grading $\Gamma$ in (1.1), there is an induced homomorphism of affine group schemes extending equation (1.2) (see [EK13, §1.4]):
$\eta_{\Gamma}\colon\mathbf{D}(G)\longrightarrow\mathbf{Aut}(\mathcal{A}),$ (1.6)
where $\mathbf{D}(G)$ is the diagonalizable group scheme represented by the group algebra $\mathbb{F}G$. Hence, for any $R$ in $\operatorname{\mathrm{Alg_{\mathbb{F}}}}$, the group $\mathbf{D}(G)(R)$ consists of the algebra homomorphisms $\mathbb{F}G\rightarrow R$, which can be identified with the group homomorphisms $G\rightarrow R^{\times}$, that is, with the characters of $G$ with values in $R^{\times}$. The homomorphism $\eta_{\Gamma}$ works as before: given a character $\chi\colon G\rightarrow R^{\times}$, the associated automorphism of $\mathcal{A}\otimes_{\mathbb{F}}R$ is given by the same formula as in (1.3):
$\tau_{\chi}(x)=\chi(g)x$
for all $x\in(\mathcal{A}\otimes_{\mathbb{F}}R)_{g}:=\mathcal{A}_{g}\otimes_{\mathbb{F}}R$.
In this new setting, the grading $\Gamma$ is recovered again from $\eta_{\Gamma}$, and actually in a very easy way. Take $R=\mathbb{F}G$, and let $\chi\in\mathbf{D}(G)(\mathbb{F}G)$ be the generic element, that is, the identity homomorphism $\mathbb{F}G\rightarrow\mathbb{F}G$, identified with the group homomorphism $G\rightarrow(\mathbb{F}G)^{\times}$ that takes any $g\in G$ to the element $g\in\mathbb{F}G$. Then for any $x\in\mathcal{A}_{g}$, $\tau_{\chi}(x\otimes 1)=x\otimes g$, and $\mathcal{A}_{g}$ consists of the ‘eigenvectors’ of ‘eigenvalue’ $g$ (computed inside $\mathcal{A}\otimes_{\mathbb{F}}\mathbb{F}G$) of $\tau_{\chi}$.
The image of $\eta_{\Gamma}$ is now contained in the diagonal group scheme of $\Gamma$ (see [EK13, p. 23]), which the reader should compare with (1.4)):
$\mathbf{Diag}(\Gamma)\colon R\longmapsto\{\varphi\in\operatorname{\mathrm{Aut}}_{\operatorname{\mathrm{Alg_{R}}}}(\mathcal{A}\otimes_{\mathbb{F}}R)\mid\forall g\in G\ \varphi|_{\mathcal{A}_{g}\otimes_{\mathbb{F}}R}\in R^{\times}\mathrm{id}\}.$
We are now ready to explain what this note is about.
The goal is to consider the affine group scheme versions, $\mathbf{Stab}(\Gamma)$ and $\mathbf{Aut}(\Gamma)$, of the stabilizer and the automorphism group of $\Gamma$. The former is straightforward, but the latter has some subtleties.
Once this is done, the scheme theoretic version of Equation (1.5) will be proven (Theorem 2.1). Again, the proof for $\mathbf{Aut}(\Gamma)$ requires extra care.
Recall that if $\mathbf{H}$ is a closed subgroup of $\mathbf{G}$, the centralizer and normalizer of $\mathbf{H}$ in $\mathbf{G}$ are the subschemes of $\mathbf{G}$ that act as follows on objects:
$\begin{split}\mathbf{Cent}_{\mathbf{G}}(\mathbf{H})\colon R&\longmapsto\{g\in\mathbf{G}(R)\mid\text{$\forall R$-algebra $R^{\prime}$ and }\forall h\in\mathbf{G}(R^{\prime})\ ghg^{-1}=h\},\\ \mathbf{Norm}_{\mathbf{G}}(\mathbf{H})\colon R&\longmapsto\{g\in\mathbf{G}(R)\mid\text{$\forall R$-algebra $R^{\prime}$}\ g\mathbf{H}(R^{\prime})g^{-1}=\mathbf{H}(R^{\prime})\}.\end{split}$
Here there is an abuse of notation, we denote by $g$ both an element in $\mathbf{G}(R)$ and its image in $\mathbf{G}(R^{\prime})$ under $\mathbf{G}(\iota)$, for $\iota\colon R\rightarrow R^{\prime}$ the algebra homomorphism determining the structure of $R^{\prime}$ as an $R$-algebra. This should create no confusion.
Once the goal above is achieved in Section 2, it is natural to consider the quotient
$\mathbf{Aut}(\Gamma)/\mathbf{Stab}(\Gamma)=\mathbf{Norm}_{\mathbf{Aut}(\mathcal{A})}\bigl(\mathbf{Diag}(\Gamma)\bigr)/\mathbf{Cent}_{\mathbf{Aut}(\mathcal{A})}\bigl(\mathbf{Diag}(\Gamma)\bigr).$
In Section 3, this scheme theoretic quotient will be shown to be a constant group scheme $\mathsf{W}(\Gamma)$ (Proposition 3.2), called the Weyl group scheme of $\Gamma$; but its associated ordinary group (the group of rational points over $\mathbb{F}$) is not, in general, our previous Weyl group $W(\Gamma)$, as shown in Example 3.3, but rather the Weyl group $W(\Gamma_{\overline{\mathbb{F}}})$ of the grading $\Gamma_{\overline{\mathbb{F}}}$ obtained on $\mathcal{A}\otimes_{\mathbb{F}}\overline{\mathbb{F}}$ by extension of scalars, where $\overline{\mathbb{F}}$ denotes an algebraic closure of $\mathbb{F}$.
In what follows $\mathcal{A}$ will always denote a finite-dimensional nonassociative algebra over an arbitrary ground field $\mathbb{F}$. Unadorned tensor products will be assumed to be over $\mathbb{F}$.
2. The automorphism group scheme of a grading
In order to define the automorphism group scheme of a grading we need a small detour to deal with gradings on vector spaces.
Given a finite-dimensional vector space $\mathcal{V}$ over a field $\mathbb{F}$ and a decomposition $\Gamma:\mathcal{V}=\bigoplus_{s\in S}\mathcal{V}_{s}$ as a direct sum of subspaces, denote by $\mathrm{GL}(\mathcal{V},\Gamma)$ the subgroup of the general linear group $\mathrm{GL}(\mathcal{V})$ consisting of those linear automorphisms $\varphi$ of $\mathcal{V}$ such that, for any $s\in\operatorname{\mathrm{Supp}}(\Gamma):=\{s\in S\mid\mathcal{V}_{s}\neq 0\}$, there is an element $\sigma(s)\in\operatorname{\mathrm{Supp}}(\Gamma)$ such that $\varphi(\mathcal{V}_{s})=\mathcal{V}_{\sigma(s)}$. Under these circumstances, $\sigma$ is a permutation of $\operatorname{\mathrm{Supp}}(\Gamma)$ ($\sigma\in\operatorname{\mathrm{Sym}}(\operatorname{\mathrm{Supp}}(\Gamma))$) with the property that $\dim_{\mathbb{F}}\mathcal{V}_{s}=\dim_{\mathbb{F}}\mathcal{V}_{\sigma(s)}$ for all $s\in\operatorname{\mathrm{Supp}}(\Gamma)$.
The affine group scheme version $\mathbf{GL}(\mathcal{V},\Gamma)$ works as follows. Any algebra $R$ in $\operatorname{\mathrm{Alg_{\mathbb{F}}}}$ without proper idempotents is sent to
$\mathbf{GL}(\mathcal{V},\Gamma)(R):=\{\varphi\in\operatorname{\mathrm{Aut}}_{R-\operatorname{\mathrm{Mod}}}(\mathcal{V}\otimes R)\mid\\ \exists\sigma\in\operatorname{\mathrm{Sym}}(\operatorname{\mathrm{Supp}}(\Gamma))\ \text{such that}\ \varphi(\mathcal{V}_{s}\otimes R)=\mathcal{V}_{\sigma(s)}\otimes R\ \forall s\in\operatorname{\mathrm{Supp}}(\Gamma)\}.$
For an arbitrary algebra $R$ in $\operatorname{\mathrm{Alg_{\mathbb{F}}}}$ this must be modified in order to get an affine group scheme:
$\begin{split}\mathbf{GL}(\mathcal{V},\Gamma)(R)&:=\{\varphi\in\operatorname{\mathrm{Aut}}_{R\textrm{-Mod}}(\mathcal{V}\otimes R)\mid\exists n\in\mathbb{N},\,\exists e_{1},\ldots,e_{n}\ \text{nonzero}\\ &\qquad\ \text{orthogonal idempotents in $R$ with}\ 1=e_{1}+\cdots+e_{n},\\ &\qquad\ \text{and }\exists\sigma_{1},\ldots,\sigma_{n}\in\operatorname{\mathrm{Sym}}(\operatorname{\mathrm{Supp}}(\Gamma)),\ \text{such that}\\ &\qquad\varphi(\mathcal{V}_{s}\otimes Re_{i})=\mathcal{V}_{\sigma_{i}(s)}\otimes Re_{i}\ \forall s\in\operatorname{\mathrm{Supp}}(\Gamma),\,\forall i=1,\ldots,n\}.\end{split}$ (2.1)
That is, $\mathbf{GL}(\mathcal{V},\Gamma)(R)$ consists of those $R$-linear automorphisms of $\mathcal{V}\otimes R$ for which there are nonzero orthogonal idempotents $e_{1},\ldots,e_{n}$, with $1=e_{1}+\cdots+e_{n}$, such that the restriction of $\varphi$ to each $\mathcal{V}\otimes Re_{i}$ permutes the nonzero components $\mathcal{V}_{s}\otimes Re_{i}$. Note that the permutation may be different for each $i=1,\ldots,n$.
For any homomorphism $f\colon R\rightarrow S$ in $\operatorname{\mathrm{Alg_{\mathbb{F}}}}$, the definition of $\mathbf{GL}(\mathcal{V},\Gamma)(f)$ is just given by extension of scalars from $R$ to $S$ via $f$.
It is not difficult to check that $\mathbf{GL}(\mathcal{V},\Gamma)$ is indeed an affine group scheme (i.e., representable). For this, let $H$ be the subgroup of $\operatorname{\mathrm{Sym}}(\operatorname{\mathrm{Supp}}(\Gamma))$ consisting of those permutations $\sigma$ with $\dim_{\mathbb{F}}\mathcal{V}_{s}=\dim_{\mathbb{F}}\mathcal{V}_{\sigma(s)}$ for any $s\in\operatorname{\mathrm{Supp}}(\Gamma)$. Write $\operatorname{\mathrm{Supp}}(\Gamma)=\{s_{1},\ldots,s_{m}\}$ and take bases $(v_{1},\ldots,v_{n_{1}})$ of $\mathcal{V}_{s_{1}}$, $(v_{n_{1}+1},\ldots,v_{n_{1}+n_{2}})$ of $\mathcal{V}_{s_{2}}$, …., $(v_{n_{1}+\cdots+n_{m-1}+1},\ldots,v_{n_{1}+\cdots+n_{m}})$ of $\mathcal{V}_{s_{m}}$. Write $I_{1}=\{1,\ldots,n_{1}\}$, $I_{2}=\{n_{1}+1,\ldots,n_{1}+n_{2}\}$, …, $I_{m}=\{n_{1}+\cdots+n_{m-1}+1,\ldots n_{1}+\cdots+n_{m}\}$. Then $\mathbf{GL}(\mathcal{V},\Gamma)$ is represented by the algebra
$\mathcal{A}=\prod_{\sigma\in H}\mathcal{A}_{\sigma},$
where, for each $\sigma\in H$, the algebra $\mathcal{A}_{\sigma}$ is the quotient of the polynomial algebra
$\mathbb{F}[X_{ij},Y\mid 1\leq i,j\leq n_{1}+\cdots+n_{m}]$
by the ideal generated by $\det(X_{ij})Y-1$ and by the $X_{ij}$’s, with $i\in I_{p}$, $j\in I_{q}$ and $s_{q}\neq\sigma(s_{p})$:
$\mathcal{A}_{\sigma}=\mathbb{F}[X_{ij},Y]/\left(\det(X_{ij})Y-1,\,X_{ij}\mid i\in I_{p},\,j\in I_{q},\,s_{q}\neq\sigma(s_{p})\right).$
Note that for $\sigma=\mathrm{id}$, the algebra $\mathcal{A}_{\mathrm{id}}$ represents the stabilizer $\mathbf{Stab}(\Gamma)$, that takes any $R$ to the $R$-linear automorphisms of $\mathcal{V}\otimes R$ fixing each component $\mathcal{V}_{s}\otimes R$.
We are now in position to define the automorphism group scheme of a grading.
Let $\Gamma:\mathcal{A}=\bigoplus_{g\in G}\mathcal{A}_{g}$ be a grading by an abelian group $G$ of our algebra $\mathcal{A}$. As before, we may consider the stabilizer affine group scheme $\mathbf{Stab}(\Gamma)$, that takes any $R$ in $\operatorname{\mathrm{Alg_{\mathbb{F}}}}$ to the group
$\mathbf{Stab}(\Gamma)(R):=\{\varphi\in\operatorname{\mathrm{Aut}}_{\operatorname{\mathrm{Alg_{R}}}}(\mathcal{A}\otimes R)\mid\varphi(\mathcal{A}_{g}\otimes R)=\mathcal{A}_{g}\otimes R\ \forall g\in G\},$
and acts naturally on homomorphisms in $\operatorname{\mathrm{Alg_{\mathbb{F}}}}$. The automorphism group scheme of the grading $\Gamma$, denoted by $\mathbf{Aut}(\Gamma)$, is defined simply as the intersection
$\mathbf{Aut}(\Gamma):=\mathbf{Aut}(\mathcal{A})\cap\mathbf{GL}(\mathcal{A},\Gamma).$ (2.2)
Note that $\mathbf{Aut}(\Gamma)(\mathbb{F})$ is the traditional group $\operatorname{\mathrm{Aut}}(\Gamma)$ of automorphisms of $\Gamma$.
The next theorem is the main result of this note. It shows that the situation for algebraically closed fields of characteristic 0 extends to arbitrary fields if affine group schemes are used.
Theorem 2.1.
Let $\Gamma:\mathcal{A}=\bigoplus_{g\in G}\mathcal{A}_{g}$ be a grading by an abelian group on a finite-dimensional algebra $\mathcal{A}$ over an arbitrary field $\mathbb{F}$. Then,
  1. (1)
    $\mathbf{Stab}(\Gamma)$ is the centralizer in $\mathbf{Aut}(\mathcal{A})$ of the diagonal group scheme $\mathbf{Diag}(\Gamma)$:
    $\mathbf{Stab}(\Gamma)=\mathbf{Cent}_{\mathbf{Aut}(\mathcal{A})}\bigl(\mathbf{Diag}(\Gamma)\bigr).$
  2. (2)
    $\mathbf{Aut}(\Gamma)$ is the normalizer in $\mathbf{Aut}(\mathcal{A})$ of the diagonal group scheme $\mathbf{Diag}(\Gamma)$:
    $\mathbf{Aut}(\Gamma)=\mathbf{Norm}_{\mathbf{Aut}(\mathcal{A})}\bigl(\mathbf{Diag}(\Gamma)\bigr).$
Proof.
For the first part, $\mathbf{Stab}(\Gamma)$ is clearly a subgroup scheme of the centralizer $\mathbf{Cent}_{\mathbf{Aut}(\mathcal{A})}\bigl(\mathbf{Diag}(\Gamma)\bigr)$.
For the converse, let $R$ be in $\operatorname{\mathrm{Alg_{\mathbb{F}}}}$ and let $\varphi$ be an element of the centralizer $\mathbf{Cent}_{\mathbf{Aut}(\mathcal{A})}\bigl(\mathbf{Diag}(\Gamma)\bigr)(R)$. Let $R^{\prime}=RG$ be the group algebra of $G$ over $R$ and consider the ‘generic’ diagonal automorphism $\psi\in\mathbf{Diag}(\Gamma)(R^{\prime})$ given by
$\psi(a_{g}\otimes 1)=a_{g}\otimes g,$
for any $g\in G$ and $a_{g}\in\mathcal{A}_{g}$. As a consequence, $\mathcal{A}_{g}\otimes R^{\prime}$ is the eigenspace of $\psi$ of eigenvalue $g\in R^{\prime}$. Then, extending $\varphi$ to $R^{\prime}$, we get, for any $g\in G$ and $a_{g}\in\mathcal{A}_{g}$,
$\psi\varphi(a_{g}\otimes 1)=\varphi\psi(a_{g}\otimes 1)=\varphi(a_{g}\otimes g)=\varphi(a_{g}\otimes 1)g,$
and this shows that $\varphi(a_{g}\otimes 1)$ lies in $\mathcal{A}_{g}\otimes R$, proving that $\varphi$ belongs to $\mathbf{Stab}(\Gamma)(R)$.
For the second part, write $\mathbf{N}=\mathbf{Norm}_{\mathbf{Aut}(\mathcal{A})}\bigl(\mathbf{Diag}(\Gamma)\bigr)$, and recall that for any $R$ in $\operatorname{\mathrm{Alg_{\mathbb{F}}}}$:
$\mathbf{N}(R):=\{\varphi\in\mathbf{Aut}(\mathcal{A})(R)\mid\varphi\mathbf{Diag}(\Gamma)(R^{\prime})\varphi^{-1}=\mathbf{Diag}(\Gamma)(R^{\prime})\ \forall R^{\prime}\text{ in}\,\operatorname{\mathrm{Alg_{R}}}\},$
while an element $\varphi\in\mathbf{Aut}(\mathcal{A})(R)$ lies in $\mathbf{Aut}(\Gamma)(R)$ if and only if there are nonzero orthogonal idempotents $e_{1},\ldots,e_{n}\in R$ with $1=e_{1}+\cdots+e_{n}$, and permutations $\sigma_{1},\ldots,\sigma_{n}$ of $\operatorname{\mathrm{Supp}}(\Gamma)$ such that
$\varphi(\mathcal{A}_{g}\otimes Re_{i})=\mathcal{A}_{\sigma_{i}(g)}\otimes Re_{i}$
for all $i=1,\ldots,n$ and all $g\in\operatorname{\mathrm{Supp}}(\Gamma)$. For any $R^{\prime}$ in $\operatorname{\mathrm{Alg_{R}}}$ and any $\delta\in\mathbf{Diag}(\Gamma)(R^{\prime})$, $\delta$ acts on $\mathcal{A}_{\sigma_{i}(g)}\otimes R^{\prime}$ by multiplication by a scalar $r^{\prime}$, and hence it acts on $\mathcal{A}_{\sigma_{i}(g)}\otimes R^{\prime}e_{i}$ by multiplication by $r^{\prime}e_{i}$. As a consequence, $\varphi^{-1}\delta\varphi$ acts on $\mathcal{A}_{g}\otimes R^{\prime}e_{i}$ by multiplication by $r^{\prime}e_{i}$. It follows then that $\varphi^{-1}\delta\varphi$ acts diagonally, and similarly for $\varphi\delta\varphi^{-1}$. Hence, $\varphi\mathbf{Diag}(\Gamma)(R^{\prime})\varphi^{-1}=\mathbf{Diag}(\Gamma)(R^{\prime})$ and $\varphi\in\mathbf{N}(R)$. We have thus proved that $\mathbf{Aut}(\Gamma)$ is a subgroup scheme of $\mathbf{N}$.
Take now an automorphism $\varphi\in\mathbf{N}(R)$. In order to check that $\varphi$ lies in $\mathbf{Aut}(\Gamma)(R)$ we will make some reductions:
  1. (1)
    We may assume that $R$ is finitely generated:
    Indeed, fix a basis $\{v_{1},\ldots,v_{n}\}$ of $\mathcal{A}$ consisting of homogeneous elements. The matrices of $\varphi$ and $\varphi^{-1}$ in the basis $\{v_{i}\otimes 1\mid 1\leq i\leq n\}$ of $\mathcal{A}\otimes R$ involve only a finite number of elements of $R$. Let $R^{\prime}$ be the subalgebra of $R$ generated by these elements. Let $\iota\colon R^{\prime}\rightarrow R$ be the inclusion. Then, our $\varphi\in\mathbf{N}(R)$ is the image under $\mathbf{N}(\iota)$ of an element in $\mathbf{N}(R^{\prime})$ and we can substitute $R$ by $R^{\prime}$.
  2. (2)
    We may assume that $R$ has no proper idempotents:
    Since we are assuming now that $R$ is finitely generated, take a family of nonzero orthogonal idempotents $e_{1},\ldots,e_{m}$ with $m$ maximal (this number is bounded by the number of irreducible components of the prime spectrum of $R$ because, for any $e=e^{2}$, the set of prime ideals containing $e$ is both open and closed). Then $1=e_{1}+\cdots+e_{m}$, as otherwise we may add the idempotent $1-(e_{1}+\cdots+e_{m})$ to our family, and hence $R=Re_{1}\oplus\cdots\oplus Re_{m}$ and each $Re_{i}$ has no proper idempotent. Now it is enough to note that $\mathbf{N}(R)=\mathbf{N}(Re_{1})\times\cdots\times\mathbf{N}(Re_{m})$.
Hence assume that $\varphi$ is an element of $\mathbf{N}(R)$ with $R$ containing no proper idempotents. As above, consider the $\mathbb{F}$-algebra $R^{\prime}=RG$ (group algebra over $R$), and the diagonal element $\psi\in\mathbf{Diag}(\Gamma)(R^{\prime})$ given by $\psi(a_{g}\otimes 1)=a_{g}\otimes g$ for any $g\in G$ and $a_{g}\in\mathcal{A}_{g}$. Since $\varphi$ is in $\mathbf{N}(R)$ we have $\varphi^{-1}\psi\varphi\in\mathbf{Diag}(\Gamma)(R^{\prime})$, so for any $g\in\operatorname{\mathrm{Supp}}(\Gamma)$ there is an invertible scalar $s_{g}\in R^{\prime}$ such that
$\varphi^{-1}\psi\varphi|_{\mathcal{A}_{g}\otimes R^{\prime}}=s_{g}\mathrm{id}.$
Thus, $s_{g}=r_{1}h_{1}+\cdots+r_{n}h_{n}$ for some $h_{1},\ldots,h_{n}\in G$ and nonzero scalars $r_{i}\in R$.
For $a\in\mathcal{A}_{g}$ we want to prove that $\varphi(a\otimes 1)$ is homogeneous and that its degree depends only on $g$. We have
$\psi\varphi(a\otimes 1)=\varphi\bigl(\varphi^{-1}\psi\varphi\bigr)(a\otimes 1)=\varphi(a\otimes s_{g})=\varphi(a\otimes 1)s_{g}.$ (2.3)
Write $\varphi(a\otimes 1)=\sum_{h\in G}a_{h}$ (finite sum), with $a_{h}\in\mathcal{A}_{h}\otimes R$. Then also
$\psi\varphi(a\otimes 1)=\sum_{h\in G}\psi(a_{h})=\sum_{h\in G}a_{h}h.$
The last equation, together with (2.3), gives
$\sum_{h\in G}a_{h}h=\sum_{h\in G}a_{h}s_{g}\ \,(\in\mathcal{A}\otimes R^{\prime}),$ (2.4)
and since $\mathcal{A}\otimes R^{\prime}=\bigoplus_{g\in G}\bigoplus_{h\in G}\mathcal{A}_{g}\otimes Rh$ we conclude that
$a_{h}=0\text{\quad for }h\neq h_{1},\ldots,h_{n},$ (2.5)
$a_{h_{i}}=a_{h_{i}}r_{i}\text{\quad for $i=1,\ldots,n$},$
$a_{h_{i}}r_{j}=0\text{\quad for $i\neq j$}.$
Let $P$ be a prime ideal of $R$. For any $b\in\mathcal{A}\otimes R$ denote by $\bar{b}$ its image in $\mathcal{A}\otimes R/P$, and for any $r\in R$ denote by $\bar{r}$ its image in the integral domain $R/P$. As $s_{g}=r_{1}h_{1}+\cdots+r_{n}h_{n}$ is invertible in $R^{\prime}=RG$, it follows that $r_{1}+\cdots+r_{n}$ is invertible in $R$, because of the counit homomorphism $RG\rightarrow R$ taking any $g\in G$ to $1$. Therefore $\bar{r}_{1}+\cdots+\bar{r}_{n}$ is invertible in $R/P$.
Also, since $\varphi$ is an automorphism, the induced map $\bar{\varphi}\in\operatorname{\mathrm{Aut}}_{\operatorname{\mathrm{Alg}}_{R/P}}(\mathcal{A}\otimes R/P)$ is also an automorphism, so $0\neq\bar{\varphi}(a\otimes 1)=\bar{a}_{h_{1}}+\cdots+\bar{a}_{h_{n}}$. In particular, there is an index $i$ with $\bar{a}_{h_{i}}\neq 0$. From (2.5) we get $\bar{a}_{h_{i}}\bar{r}_{i}=\bar{a}_{h_{i}}$, so $\bar{r}_{i}\neq 0$, and since $R/P$ is an integral domain, its action on the free $R/P$-module $\mathcal{A}\otimes R/P$ has no torsion, so $\bar{r}_{j}=0$, because $\bar{a}_{h_{i}}\bar{r}_{j}=0$ by (2.5).
Hence, for any prime ideal $P$ of $R$ there is a unique index $i$, $1\leq i\leq n$, such that $\bar{r}_{i}\neq 0$, that is, $r_{i}\not\in P$. We thus get a locally constant map
$\Phi\colon\operatorname{Spec}(R)\longrightarrow\{1,\ldots,n\},$
where we consider the discrete topology on the right, with
$\Phi^{-1}(i)=\{P\in\operatorname{Spec}(R)\mid r_{i}\not\in P\}$
which is the basic open set $D(r_{i})$ in $\operatorname{Spec}(R)$.
But $\operatorname{Spec}(R)$ is connected because $R$ does not contain proper idempotents, so we conclude that there is an index $i$ which, after relabeling, we may assume to be $1$, such that for any prime ideal $P$, $r_{1}\not\in P$ while $r_{2},\ldots,r_{m}\in P$. It follows that $r_{1}$ is a unit in $R$, as it is not contained in any maximal ideal, while $r_{2},\ldots,r_{m}$ are nilpotent. Now from (2.5) we get $a_{h_{j}}r_{1}=0$ for $j\neq 1$, and hence $a_{h_{j}}=0$ for $j\neq 1$ and the element $s_{g}$ in (2.3) is just $r_{1}h_{1}$. We conclude that $\varphi(a\otimes 1)$ belongs to $\mathcal{A}_{h_{1}}\otimes R$, and $h_{1}$ depends only on $g\in\operatorname{\mathrm{Supp}}(\Gamma)$ and not on the element $a\in\mathcal{A}_{g}$.
Therefore, we have shown that for any $g\in\operatorname{\mathrm{Supp}}(\Gamma)$, there is an element $h\in\operatorname{\mathrm{Supp}}(\Gamma)$ with $\varphi(\mathcal{A}_{g}\otimes R)\subseteq\mathcal{A}_{h}\otimes R$, and the same works for $\varphi^{-1}$. In conclusion, $\varphi$ permutes the homogeneous components, so $\varphi\in\mathbf{Aut}(\Gamma)(R)$, as required. ∎
A remark is in order.
Remark 2.2.
Let $\Gamma:\mathcal{A}=\bigoplus_{g\in G}\mathcal{A}_{g}$ be a grading by an abelian group on a finite-dimensional algebra $\mathcal{A}$ over an arbitrary field $\mathbb{F}$ as in Theorem 2.1, and let $\eta_{\Gamma}\colon\mathbf{D}(G)\rightarrow\mathbf{Aut}(\mathcal{A})$ be the associated morphism of group schemes in (1.6). If $H$ is the subgroup of $G$ generated by $\operatorname{\mathrm{Supp}}(\Gamma)$, then $\eta_{\Gamma}$ factors naturally as
$\mathbf{D}(G)$$\mathbf{D}(H)$$\mathbf{Aut}(\mathcal{A}),$}
where the first morphism is the restriction to $H$ (a quotient map) and the second gives a closed imbedding. We can then identify $\mathbf{D}(H)$ with a subgroup of $\mathbf{Aut}(\mathcal{A})$. It is the image of the morphism $\eta_{\Gamma}$.
The proof of the first part of Theorem 2.1 works substituting $\mathbf{Diag}(\Gamma)$ by $\mathbf{D}(H)$, and hence it gives
$\mathbf{Stab}(\Gamma)=\mathbf{Cent}_{\mathbf{Aut}(\mathcal{A})}\bigl(\mathbf{D}(H)\bigr).$
For the second part, the proof of the converse works too, obtaining
$\mathbf{Aut}(\Gamma)\geq\mathbf{Norm}_{\mathbf{Aut}(\mathcal{A})}\bigl((\mathbf{D}(H)\bigr),$
but the equality fails in general, as Example 2.3 shows.
Example 2.3.
Let $\mathcal{A}=\mathbb{F}u\oplus\mathbb{F}v$ be a two-dimensional algebra with trivial multiplication: $u^{2}=uv=vu=v^{2}=0$, graded by the cyclic group $G=\langle g\rangle$ of order $6$, with the elements $u$ and $v$ homogeneous of degrees $\deg(u)=g^{2}$ and $\deg(v)=g^{3}$. The group $G$ is generated by the support $\{g^{2},g^{3}\}$ of this grading $\Gamma$.
For any algebra $R$ in $\operatorname{\mathrm{Alg_{\mathbb{F}}}}$, the elements of $\mathbf{D}(G)(R)$, considered as elements in $\mathbf{Aut}(\mathcal{A})(R)$ are the automorphisms
$\tau_{\chi}\colon u\mapsto\chi(g^{2})u,\ v\mapsto\chi(g^{3})v$
for a character (i.e., group homomorphism) $\chi\colon G\rightarrow R^{\times}$.
The automorphism $\varphi$ that swaps $u$ and $v$ clearly belongs to $\mathbf{Aut}(\Gamma)(\mathbb{F})=\operatorname{\mathrm{Aut}}(\Gamma)$, but if $\chi\colon G\rightarrow(\mathbb{F}G)^{\times}$ is the character with $\chi(g)=g$, then there is no character $\xi\colon G\rightarrow(\mathbb{F}G)^{\times}$ with $\varphi\tau_{\chi}\varphi^{-1}=\tau_{\xi}$, because this would force $\xi(g^{2})=g^{3}$, $\xi(g^{3})=g^{2}$, which is impossible, as we would have $\xi(g^{5})=g^{5}$, and $\xi=\chi$ because $g^{5}$ is a generator of $G$, a contradiction.
Therefore, $\varphi$ lies in $\mathbf{Aut}(\Gamma)(\mathbb{F})$ but not in $\mathbf{Norm}_{\mathbf{Aut}(\mathcal{A})}\bigl(\mathbf{D}(G)\bigr)(\mathbb{F})$.
The computation of $\mathbf{Diag}(\Gamma)$ and of $\mathbf{Aut}(\Gamma)$ is left as an exercise for the reader.
Let us finish this section with a useful remark.
Remark 2.4.
Assume that the ground field $\mathbb{F}$ is algebraically closed and that $\mathbf{H}$ is a smooth closed subgroup scheme of the affine algebraic group scheme $\mathbf{G}$. Then
$\mathbf{Norm}_{\mathbf{G}}(\mathbf{H})(\mathbb{F})=\operatorname{\mathrm{Norm}}_{\mathbf{G}(\mathbb{F})}\bigl(\mathbf{H}(\mathbb{F})\bigr).$ (2.6)
In particular, over an algebraically closed field of characteristic 0, so that all affine group schemes are smooth, if $\Gamma:\mathcal{A}=\bigoplus_{g\in G}\mathcal{A}_{g}$ is a grading by an abelian group on a a finite-dimensional algebra $\mathcal{A}$, then we recover the result [EK13, (1.6)] that $\operatorname{\mathrm{Aut}}(\Gamma)=\mathbf{Aut}(\Gamma)(\mathbb{F})$ is $\operatorname{\mathrm{Norm}}_{\operatorname{\mathrm{Aut}}(\mathcal{A})}\bigl(\operatorname{\mathrm{Diag}}(\Gamma)\bigr)=\mathbf{Norm}_{\mathbf{Aut}(\mathcal{A})(\mathbb{F})}\bigl(\mathbf{Diag}(\Gamma)(\mathbb{F})\bigr)$.
To prove (2.6), let $\mathcal{H}$ be the Hopf algebra representing $\mathbf{G}$, and let $\mathcal{I}$ be the Hopf ideal defining $\mathbf{H}$. As this is smooth, $\mathcal{I}=\bigcap_{x\in\mathbf{H}(\mathbb{F})}\mathcal{M}_{x}$, where $\mathcal{M}_{x}$ is the maximal ideal corresponding to $x$: $\mathcal{M}_{x}=\ker x\colon\mathcal{H}\rightarrow\mathbb{F}$.
Let $\Delta$ be the comultiplication on $\mathcal{H}$, $S$ its antipode, and use Sweedler’s notation to write $(\Delta\otimes\mathrm{id})\Delta(a)=\sum a_{(1)}\otimes a_{(2)}\otimes a_{(3)}$. Any $g\in\mathbf{G}(\mathbb{F})=\operatorname{\mathrm{Hom}}_{\operatorname{\mathrm{Alg_{\mathbb{F}}}}}(\mathcal{H},\mathbb{F})$ acts by conjugation on $\mathcal{H}$:
$\iota_{g}(a)=\sum g(a_{(1)})gS(a_{(3)})a_{(2)},$
so for any $R$ in $\operatorname{\mathrm{Alg_{\mathbb{F}}}}$ and $h\in\mathbf{G}(R)$:
$ghg^{-1}\colon a\mapsto\sum g(a_{(1)})h(a_{(2)})gS(a_{(3)})=h(\iota_{g}(a)).$
Hence $ghg^{-1}=h\iota_{g}$. Note that $\iota_{g_{1}g_{2}}=\iota_{g_{2}}\iota_{g_{1}}$ for any $g_{1},g_{2}\in\mathbf{G}(\mathbb{F})$.
For $g\in\mathbf{G}(\mathbb{F})$, $\iota_{g}(\mathcal{I})=\mathcal{I}$ if and only if for all $R\in\operatorname{\mathrm{Alg_{\mathbb{F}}}}$ and all $h\in\mathbf{G}(R)$ we have that $h(\mathcal{I})=0$ if and only if $ghg^{-1}(\mathcal{I})=0$, which holds if and only if $g\in\mathbf{Norm}_{\mathbf{G}}(\mathbf{H})(\mathbb{F})$.
Note that $\mathbf{Norm}_{\mathbf{G}}(\mathbf{H})(\mathbb{F})$ is contained in $\operatorname{\mathrm{Norm}}_{\mathbf{G}(\mathbb{F})}\bigl(\mathbf{H}(\mathbb{F})\bigr)$ trivially. On the other hand, for $g\in\operatorname{\mathrm{Norm}}_{\mathbf{G}(\mathbb{F})}\bigl(\mathbf{H}(\mathbb{F})\bigr)$, $ghg^{-1}\in\mathbf{H}(\mathbb{F})$ for any $h\in\mathbf{H}(\mathbb{F})$, and hence $h\bigl(\iota_{g}(\mathcal{I})\bigr)=0$ for any $h\in\mathbf{H}(\mathbb{F})$, so that $\iota_{g}(\mathcal{I})$ is contained in $\mathcal{M}_{x}$ for all $x\in\mathbf{H}(\mathbb{F})$, and we conclude that $\iota_{g}(\mathcal{I})$ is contained in $\mathcal{I}$. As $g^{-1}$ lies in $\operatorname{\mathrm{Norm}}_{\mathbf{G}(\mathbb{F})}\bigl(\mathbf{H}(\mathbb{F})\bigr)$ too, we get also $\iota_{g^{-1}}(\mathcal{I})\subseteq\mathcal{I}$, so $\iota_{g}(\mathcal{I})=\mathcal{I}$ and $g$ is in $\mathbf{Norm}_{\mathbf{G}}(\mathbf{H})(\mathbb{F})$, as required.
If $\mathbf{H}$ is not smooth the equality $\mathbf{Norm}_{\mathbf{G}}(\mathbf{H})(\mathbb{F})=\operatorname{\mathrm{Norm}}_{\mathbf{G}(\mathbb{F})}\bigl(\mathbf{H}(\mathbb{F})\bigr)$ is not valid in general.
3. Weyl groups
This last section is devoted to define the Weyl group scheme of a grading as a constant group scheme.
Given a grading $\Gamma:\mathcal{A}=\bigoplus_{g\in G}\mathcal{A}_{g}$ by an abelian group on a a finite-dimensional algebra $\mathcal{A}$ over an arbitrary field $\mathbb{F}$, the definition of $\mathbf{Aut}(\Gamma)$ gives at once a natural morphism of affine group schemes
$\mathbf{Aut}(\Gamma)\rightarrow\mathsf{Sym}\bigl(\operatorname{\mathrm{Supp}}(\Gamma)\bigr),$ (3.1)
with kernel $\mathbf{Stab}(\Gamma)$, and where the scheme on the right is the constant group scheme associated to the symmetric group on $\operatorname{\mathrm{Supp}}(\Gamma)$. (Recall that given a finite group $G$, its associated constant group scheme $\mathsf{G}$ assigns to any $R$ in $\operatorname{\mathrm{Alg_{\mathbb{F}}}}$ without proper idempotents the group $G$.)
Definition 3.1.
The image of the morphism in (3.1) is called the Weyl group scheme of the grading $\Gamma$. It will be denoted by $\mathsf{W}(\Gamma)$.
Since any subgroup of a constant affine group scheme is constant, $\mathsf{W}(\Gamma)$ is a constant group scheme, and there appears the short exact sequence
[…]
If $\overline{\mathbb{F}}$ denotes an algebraic closure of $\mathbb{F}$, since $\mathbf{Aut}(\Gamma)\twoheadrightarrow\mathsf{W}(\Gamma)$ is a quotient map, the induced group homomorphism $\mathbf{Aut}(\Gamma)(\overline{\mathbb{F}})\twoheadrightarrow\mathsf{W}(\Gamma)(\overline{\mathbb{F}})$ is surjective (see, e.g., [EK13, Theorem A.48]), and we get the short exact sequence of groups
[…]
Denote by $\Gamma_{\overline{\mathbb{F}}}$ the $G$-grading induced by $\Gamma$ on $\mathcal{A}_{\overline{\mathbb{F}}}=\mathcal{A}\otimes\overline{\mathbb{F}}$: $(\mathcal{A}_{\overline{\mathbb{F}}})_{g}=\mathcal{A}_{g}\otimes\overline{\mathbb{F}}$ for all $g\in G$. The last short exact sequence becomes
[…]
and this shows that $\mathsf{W}(\Gamma)(\overline{\mathbb{F}})$ is the ordinary Weyl group $W(\Gamma_{\overline{\mathbb{F}}})$. In other words:
Proposition 3.2.
The Weyl group scheme of $\Gamma$ is the constant group scheme associated to the Weyl group of $\Gamma_{\overline{\mathbb{F}}}$.
In general, $\mathsf{W}(\Gamma)$ is not the constant group scheme associated to the ordinary Weyl group $W(\Gamma)$, as the following example shows.
Example 3.3.
Let $\mathcal{A}$ be the associative $\mathbb{Q}$-algebra $\mathcal{A}=\mathbb{Q}1\oplus\mathbb{Q}u\oplus\mathbb{Q}u^{2}$, with multiplication determined by $u^{3}=2$. That is, $\mathcal{A}$ is, up to isomorphism, the field $\mathbb{Q}(\sqrt[3]{2})$. The direct sum above gives a grading $\Gamma$ on $\mathcal{A}$ by the cyclic group $C_{3}=\langle g\rangle$ of order $3$, where $\deg(1)=e$, $\deg(u)=g$ and $\deg(u^{2})=g^{2}$. The automorphism group of $\mathcal{A}$ is trivial, and hence so are $\operatorname{\mathrm{Stab}}(\Gamma)$, $\operatorname{\mathrm{Aut}}(\Gamma)$, and $W(\Gamma)$.
[…]
Trivially, $\mathbf{Stab}(\Gamma)$ is, up to isomorphism, the affine group scheme of third roots of unity $\boldsymbol{\mu}_{3}$, and the short exact sequence (3.2) becomes
$1\longrightarrow\boldsymbol{\mu}_{3}\longrightarrow\mathbf{Aut}(\Gamma)\longrightarrow\mathsf{C}_{2}\longrightarrow 1.$ (3.3)
Since $\mathbf{Aut}(\Gamma)(\mathbb{Q})$ is trivial, it turns out that the quotient map $\mathbf{Aut}(\Gamma)\rightarrow\mathsf{C}_{2}$ is not surjective for rational points. It also follows that the short exact sequence (3.3) does not split, because the group $\operatorname{\mathrm{Aut}}(\Gamma)=\mathbf{Aut}(\Gamma)(\mathbb{Q})$ is trivial, so it does not contain cyclic subgroups of order $2$.
Note that since $\mathcal{A}$ is an étale algebra, $\mathbf{Aut}(\mathcal{A})$ is an étale group scheme (see e.g., [KMRT98, Exercise VI.1]). Also, $\mathbf{Aut}(\mathcal{A})(\overline{\mathbb{Q}})=\operatorname{\mathrm{Aut}}(\mathcal{A}_{\overline{\mathbb{Q}}})\simeq\operatorname{\mathrm{Aut}}(\overline{\mathbb{Q}}^{3})$, which is the symmetric group of degree $3$ (and order $6$). From here we get that $\mathbf{Aut}(\Gamma)$ is the whole $\mathbf{Aut}(\mathcal{A})$.
The last result is a consequence of Theorem 2.1. It deals with fine gradings, which are those gradings $\Gamma:\mathcal{A}=\bigoplus_{g\in G}\mathcal{A}_{g}$ such that there exists no other grading $\Gamma^{\prime}:\mathcal{A}=\bigoplus_{h\in H}\mathcal{A}^{\prime}_{h}$ by an abelian group with the property that every nonzero homogeneous component $\mathcal{A}^{\prime}_{h}$ of $\Gamma^{\prime}$ is contained in a homogeneous component $\mathcal{A}_{g}$ of $\Gamma$, with at least one of these containments being strict. In a way, fine gradings are those whose homogeneous components are small, and this translates into the fact that its diagonal group scheme $\mathbf{Diag}(\Gamma)$ is a maximal diagonalizable subgroup scheme (or quasitorus) of $\mathbf{Aut}(\mathcal{A})$. (See [EK13, Proposition 1.37].)
Given a grading $\Gamma:\mathcal{A}=\bigoplus_{g\in G}\mathcal{A}_{g}$, the diagonal group scheme $\mathbf{Diag}(\Gamma)$ is a diagonalizable group scheme, and hence it is isomorphic to $\mathbf{D}(U)$ for an abelian group $U$. This is the universal group of $\Gamma$. The associated morphism $\eta_{\Gamma}\colon\mathbf{D}(G)\rightarrow\mathbf{Aut}(\mathcal{A})$ factors as $\mathbf{D}(G)\rightarrow\mathbf{D}(U)\simeq\mathbf{Diag}(\Gamma)\hookrightarrow\mathbf{Aut}(\mathcal{A})$, and this induces a group homomorphism $U\rightarrow G$. The grading $\Gamma$ can be realized as a grading by $U$, and as such, the homomorphism $U\rightarrow G$ takes bijectively the support of $\Gamma$ as an $U$-grading with the support of $\Gamma$ as a $G$-grading. Therefore, when considered as a grading by $U$, the morphism $\eta_{\Gamma}\colon\mathbf{D}(U)\rightarrow\mathbf{Aut}(\mathcal{A})$ takes $\mathbf{D}(U)$ isomorphically onto $\mathbf{Diag}(\Gamma)$.
Corollary 3.4.
Let $\mathcal{A}$ and $\mathcal{B}$ be two finite-dimensional algebras over a field $\mathbb{F}$ and let $\phi\colon\mathbf{Aut}(\mathcal{A})\rightarrow\mathbf{Aut}(\mathcal{B})$ be an isomorphism. Let $\Gamma$ be a fine grading on $\mathcal{A}$ with universal group $U$. Let $\eta_{\Gamma}\colon\mathbf{D}(U)\rightarrow\mathbf{Aut}(\mathcal{A})$ be the associated morphism, and let $\Gamma^{\prime}$ be the grading by $U$ on $\mathcal{B}$ associated to the morphism $\phi\eta_{\Gamma}$. Then $\Gamma^{\prime}$ is fine and $\phi$ restricts to isomorphisms $\mathbf{Stab}(\Gamma)\simeq\mathbf{Stab}(\Gamma^{\prime})$ and $\mathbf{Aut}(\Gamma)\simeq\mathbf{Aut}(\Gamma^{\prime})$. As a consequence, $\phi$ induces an isomorphism $\mathsf{W}(\Gamma)\simeq\mathsf{W}(\Gamma^{\prime})$.
Proof.
Since $\Gamma$ is fine, $\mathbf{Diag}(\Gamma)$ is a maximal quasitorus of $\mathbf{Aut}(\mathcal{A})$. Note that $\mathbf{Diag}(\Gamma)$ is the image under $\eta_{\Gamma}$ of $\mathbf{D}(U)$. The image under $\phi$ of $\mathbf{Diag}(\Gamma)$ is then a maximal quasitorus of $\mathbf{Aut}(\mathcal{B})$. It follows that $\Gamma^{\prime}$ is fine with universal group $U$, and that the image of $\mathbf{Diag}(\Gamma)$ under $\phi$, that is, the the image of $\mathbf{D}(U)$ under $\eta_{\Gamma}^{\prime}=\phi\eta_{\Gamma}$, is $\mathbf{Diag}(\Gamma^{\prime})$. Now Theorem 2.1 implies at once that $\phi$ restricts to isomorphisms $\mathbf{Stab}(\Gamma)\simeq\mathbf{Stab}(\Gamma^{\prime})$ and $\mathbf{Aut}(\Gamma)\simeq\mathbf{Aut}(\Gamma^{\prime})$ and, as a consequence, it induces an isomorphism $\mathsf{W}(\Gamma)\simeq\mathsf{W}(\Gamma^{\prime})$. ∎